### 08 Mar 2026 v1

# Fusion Complexity Inversion: Why Simpler Cross View Modules Outperform SSMs and Cross View Attention Transformers for Pasture Biomass Regression

Abstract
Accurate estimation of pasture biomass from agricultural imagery is critical for sustainable livestock management, yet existing methods are limited by the small, imbalanced, and sparsely annotated datasets typical of real-world monitoring. In this study, adaptation of vision foundation models to agricultural regression is systematically evaluated on the CSIRO Pasture Biomass benchmark, a 357 image dual view dataset with laboratory validated, component wise ground truth for five biomass targets, through 17 configurations spanning four backbones (EfficientNet-B3 to DINOv3-ViT-L), five cross view fusion mechanisms, and a metadata factorial. A counterintuitive principle, termed fusion complexity inversion, is uncovered: on scarce agricultural data, a two layer gated depthwise convolution () outperforms cross view attention transformers (), bidirectional SSMs (), and full Mamba (, below the no-fusion baseline). Backbone pretraining scale is found to monotonically dominate all architectural choices, with the DINOv2DINOv3 upgrade alone yielding points. Training only metadata (species, state, and NDVI) is shown to create a universal ceiling at , collapsing an 8.4 point fusion spread to 0.1 points. Actionable guidelines for sparse agricultural benchmarks are established: backbone quality should be prioritized over fusion complexity, local modules preferred over global alternatives, and features unavailable at inference excluded.
Keywords: pasture biomass estimation, foundation models, agricultural imagery, dual view fusion, sparse annotations, precision agriculture.
I Introduction
Scalable, non-destructive estimation of crop and pasture properties from imagery alone is increasingly enabled by computer vision [20]. Pasture biomass estimation, the prediction of dry matter weight of vegetation components from field photographs, is a canonical agricultural vision task requiring fine grained pattern recognition, robustness to sparse and imbalanced annotations, and generalization across geographic and seasonal conditions. Traditional methods (like rising plate meters, and destructive harvesting) cannot scale to the millions of hectares under pastoral management, motivating vision based alternatives.
Self supervised vision transformers pretrained at scale, notably DINOv2 [17] and DINOv3 [18] (up to 1.7B images), now provide general purpose encoders that transfer to narrow agricultural domains with minimal task specific data [28, 4]. Scale dependent transfer has been further demonstrated by vision language [27] and large language models [29]. A critical question remains, however: Given a powerful pretrained backbone, how much task specific complexity should be added when training data is scarce?
The CSIRO Pasture Biomass dataset [15], the first publicly available, multi-site pasture resource combining visual, spectral, and structural modalities with laboratory validated, component wise ground truth, is adopted as the benchmark. Unlike prior datasets relying on visual estimation or single site collection, CSIRO spans 19 sites across four Australian states over three years (2014–2017), with consumer grade cameras under natural lighting. Each of 357 dual view training photographs is paired with destructive cut-dry-weigh measurements: vegetation within a 70 cm 30 cm quadrat is harvested, sorted into green, dead, and clover fractions, oven dried at 70∘C for 48 h, and laboratory weighed, producing ground truth unmatched by any comparable benchmark. Three targets show significant zero inflation (up to 37.8% for clover), and right skewed distributions. Auxiliary metadata (species, state, NDVI, height, and date), available only during training, creates a realistic modality shift scenario common in agricultural deployment.
A systematic study spanning 17 configurations is presented along three axes: (1) cross view fusion complexity (identity, gated depthwise convolution, cross-view gated attention transformer, bidirectional Mamba SSM, and full Mamba SSM), (2) backbone scale (EfficientNet-B3 [30] through DINOv3-ViT-L, spanning ImageNet-1K [38] to LVD-1.7B), and (3) training only metadata fusion. All experiments are conducted with identical recipes, 5 fold stratified group cross validation, and a single 8 GB consumer GPU.
Three principal findings are established. (1) Fusion complexity inversion: a two layer gated depthwise convolution () outperforms all global alternatives, and full Mamba () falls below the no fusion baseline. (2) Foundation model scale dominance: increases monotonically from EfficientNet-B3 () to DINOv3-ViT-L (), and the DINOv2DINOv3 upgrade alone yields points. (3) Metadata fusion can harm: training only metadata collapses all fusion types to , destroying the best model’s 7.4 point advantage. Feature space analysis of image derived color indices and sensor metadata reveals moderate correlations between simple hand crafted features and biomass components, establishing an interpretable baseline that learned representations must surpass.
II Related Work
Computer vision for agricultural imagery. Crop and pasture analysis has progressed from hand crafted vegetation indices [2, 9, 20] to CNN based proximal sensing [10, 19], composition classification [33], and aerial monitoring [39]. A persistent bottleneck in agricultural vision is the scarcity of annotated data. The CSIRO Pasture Biomass dataset [15] exemplifies this: laboratory validated, component wise ground truth (destructive cut-dry-weigh) is paired with heterogeneous inputs (imagery, NDVI, and compressed height) across 19 sites in four Australian states, yet only 357 training images are provided, with significant zero inflation and right skewed targets.
Foundation models in agriculture. Self supervised pre-training at scale has been catalyzed by the vision transformer paradigm [26, 31]. DINO [4], DINOv2 [17] (LVD-142M), and DINOv3 [18] (LVD-1.7B) learn label free features adopted for remote sensing [5] and plant phenotyping [11], and complementary signals are provided by masked autoencoders [28]. Scale dependent transfer has been confirmed by vision language [27] and large language models [29], yet systematic guidance on task specific complexity atop foundation models for scarce agricultural data is lacking. This gap is addressed here by quantifying backbone–fusion interactions on a small agricultural benchmark.
Efficient training on sparse agricultural data. Differential learning rates [13], gradient checkpointing, mixed precision training [14], data augmentation [30, 37], and robust loss functions [36] enable fine tuning of large backbones on consumer hardware. In this study, is achieved on 357 images without external data through these strategies combined with an appropriate foundation model.
Data fusion and auxiliary metadata integration. Dual branch architectures employ cross attention [21, 22], concatenation [3, 6], or late fusion [7], and ModDrop [40] provides robustness to missing modalities. Metadata fusion is common in remote sensing [1, 8], and Gaussian process regression [34] provides a probabilistic framework for sparse data regression. Selective SSMs (Mamba [12]) have been adapted for vision through VMamba [16], with bidirectional [23, 24] and hybrid [25] variants. Five fusion paradigms are benchmarked here, and local fusion is found to dominate all global alternatives. Training only metadata is shown to act as a harmful shortcut ( points).
III Method
III-A Problem Formulation
Given a dual view pasture photograph, the task is to predict five biomass targets: Dry Green (), Dry Dead (), Dry Clover (), Green Dry Matter (GDM = Green + Clover), and Dry Total (Total = GDM + Dead). All targets are transformed to stabilize variance across right skewed distributions. The primary metric is a weighted :
where the weights reflect agronomic priorities: Dry Total (50%) is the primary indicator of carrying capacity, GDM (20%) measures digestible fraction, and the three components each receive 10%.
III-B Dual View Input Pipeline
Each input image ( pixels) captures a 70 cm 30 cm pasture quadrat from above, split vertically into left/right halves and resized to with area-based resampling (INTER_AREA). Both halves are normalized with ImageNet statistics and pass through a weight tied backbone, halving the parameter count relative to independent encoders while providing complementary spatial coverage of the quadrat. The resulting token sequences ( each) are concatenated to form a joint representation before fusion.
III-C Architecture
The model comprises four components: a weight tied backbone, a local fusion module, adaptive pooling, and compositional prediction heads (Figure˜1).
Backbone. DINOv3-ViT-L (303.08M parameters, 24 transformer layers) pretrained on LVD-1.7B is used. Each view yields a token grid of dimension 1024. Gradient checkpointing fits the dual-view forward pass within 8 GB VRAM. The backbone is fine-tuned at , while task specific layers use .
Fusion: Gated Depthwise Convolution. The concatenated token sequence is processed by two stacked GatedDepthwiseConvBlocks. Each block applies LayerNorm, sigmoid gating, depthwise 1D convolution (), linear projection with dropout (), and a residual connection:
Two stacked blocks have an effective receptive field of 9 tokens. This local operation does not attend across the full sequence: the backbone’s self attention already captures global dependencies within each view. Total fusion parameters: 4.21M.
Prediction Heads. Three independent heads (Green, Dead, and Clover) map the average pooled 1024-d vector to scalars through a two layer MLP with Softplus output:
Composite targets are computed by summation: , . Total task specific parameters: 5.79M (1.9% of the 308.87M total), comprising 4.21M fusion and 1.58M heads. No metadata is used at training or inference.
III-D Baseline Fusion Mechanisms
Four alternative fusion modules are benchmarked with the same DINOv3-ViT-L backbone, training recipe, and cross validation protocol.
Cross View Gated Attention (CVGA). The concatenated sequence is split into left and right halves, and bidirectional cross attention (8 heads, ) with sigmoid gating enables global cross view interaction at cost. Two blocks: 10.50M parameters.
Bidirectional Mamba SSM (BidirMamba). Combines local depthwise convolution () with weight tied bidirectional Mamba SSM (, expand2) for global sequence modeling. Requires FP32 due to CUDA kernel constraints. Two blocks: 17.55M parameters.
Full Mamba SSM (MambaFusionBlock). Unidirectional Mamba with expand2 and no gating or depthwise convolution overhead, making each block leaner than BidirMamba despite the same expand factor. Two blocks: 13.34M parameters.
Identity (no fusion). The concatenated sequence passes directly to pooling with no learned cross view interaction (zero parameters).
Table˜I summarizes all fusion blocks.
| Property | GatedDWConv | CVGA | BidirMamba | FullMamba | Identity |
| Receptive field | Local () | Global | Local+Global | Global | None |
| Params/block | 2.11M | 5.25M | 8.77M | 6.67M | 0 |
| Total (2 blocks) | 4.21M | 10.50M | 17.55M | 13.34M | 0 |
| AMP compatible | Yes | Yes | No | No | Yes |
III-E Metadata Injection (Ablation Variant)
For metadata inclusive experiments, the CSIRO training set’s auxiliary information is encoded into a 23 dimensional vector: State (one hot, 4d), Species (one hot, 15d), NDVI (1d), Height (1d), and cyclical sampling month (2d). A two layer MLP (, 1.12M parameters) produces a metadata embedding that is concatenated with the 1024-d pooled image features and projected back to 1024 dimensions. During training, the metadata vector is zeroed with probability per sample, and at test time, metadata is absent entirely. As shown in Section˜IV-F, this dropout rate is insufficient to prevent the metadata shortcut.
IV Experiments
IV-A Dataset
The competition subset of the CSIRO Pasture Biomass dataset [15] comprises 357 dual view photographs from 19 sites across four Australian states (2014–2017), selected from 3,187 samples through rigorous quality control. At each site, vegetation within a 70 cm 30 cm quadrat was photographed with consumer grade cameras under natural lighting, then harvested, sorted into green, dead, and clover fractions (30 g subsample), oven dried at 70∘C for 48 h, and laboratory weighed. CSIRO is the first public pasture benchmark to provide separate dead matter annotations, and to combine visual, spectral (GreenSeeker NDVI, 100 reading average), and structural (falling plate meter) modalities.
All five biomass targets are right skewed (skewness 1.4–2.8), with Dry Clover showing 37.8% zero values (Figure˜2). The heavy right tails motivate the transform, and composite targets (GDM, Dry Total) have no zeros by construction. Table˜II provides summary statistics.
| Target | Mean | Std | Min | Med. | Max | Skew | Zero% | |
| Dry Green | 0.1 | 26.6 | 25.4 | 0.0 | 20.8 | 158.0 | 1.75 | 5.0 |
| Dry Dead | 0.1 | 12.0 | 12.4 | 0.0 | 8.0 | 83.8 | 1.76 | 11.2 |
| Dry Clover | 0.1 | 6.7 | 12.1 | 0.0 | 1.4 | 71.8 | 2.84 | 37.8 |
| GDM | 0.2 | 33.3 | 24.9 | 1.0 | 27.1 | 158.0 | 1.56 | 0.0 |
| Dry Total | 0.5 | 45.3 | 28.0 | 1.0 | 40.3 | 185.7 | 1.43 | 0.0 |
Correlation structure. Figure˜3 presents the Pearson correlation matrix among all five biomass targets and four metadata variables. Dry Green and GDM are highly correlated () by construction, as are GDM and Dry Total (). Auxiliary metadata includes NDVI ( with green biomass), compressed height ( with total biomass), pasture species, state, and sampling date, all absent at test time.
Figure˜4 visualizes metadata–biomass relationships: both correlations show substantial scatter, confirming genuine but insufficient signal, and incorporating metadata degrades the best model by 7.4 points (Section˜IV-F).
Geographic and seasonal variation. Victoria and Tasmania show higher median biomass (50–55 g) than NSW and WA (30–40 g), though substantial interquartile overlap indicates state alone is a weak predictor (Figure˜5). Collection peaks in spring and autumn (Figure˜6), and species composition spans 15 types dominated by ryegrass-clover mixtures (Figure˜7), and lucerne has higher average biomass, but wide per species ranges motivate the visual only approach.
IV-B Cross-Validation Protocol
All models are evaluated with 5-fold Stratified Group K-Fold cross validation (scikit-learn, and seed 17). Folds are stratified on Dry Total quintiles, and grouped by image ID to prevent leakage. Each fold contains validation and training images. Models train for up to 50 epochs with early stopping (patience 10) on validation weighted . All 17 configurations converged before the 50 epoch maximum, with training halted by early stopping in every case. A single master seed controls all randomness across the 17 configurations.
IV-C Training Configuration
All experiments use AdamW [13] with differential learning rates ( backbone, heads), weight decay , cosine annealing with 5 epoch linear warmup, Huber loss [36] () on targets, and gradient clipping at 1.0. Differential rates preserve pretrained backbone representations while accelerating randomly initialized heads. Mixed precision (FP16) is used for AMP compatible blocks (GatedDWConv, CVGA), while Mamba blocks require FP32, consuming more VRAM. Gradient checkpointing trades computation for VRAM reduction, essential for dual view DINOv3-ViT-L on 8 GB. Augmentations (flip, rotation , shift-scale-rotate, and color jitter) are applied identically to both views through albumentations, and no test time augmentation (TTA) is used. All experiments run on a single NVIDIA RTX 4060 Laptop GPU (8 GB VRAM), with each 5 fold CV run requiring 4–8 hours.
IV-D Main Results
Table˜III presents all 17 configurations. The proposed model (DINOv3-ViT-L + 2 GatedDWConv, no metadata) achieves , outperforming all alternatives by at least 5 points. A dense cluster of DINOv3 based models occupies the – range regardless of fusion mechanism, while VMamba based models (), and EfficientNet-B3 () form progressively weaker tiers. Figure˜8(b) reveals a near linear relationship between log pretraining scale and downstream .
The median predictor (B1, ) and zero shot DINOv2+ConvNeXt (B3, ) confirm that neither constant prediction nor off the shelf features are viable, with B3’s negative score underscoring the necessity of task specific training. Among fine tuned models, a clear three tier hierarchy emerges: DINOv3 (–), VMamba (–), and EfficientNet-B3 (), aligning with pretraining data scale.
| ID | Model | Backbone | Fusion | Meta | Std | |
| B1 | Median Predictor | None | None | No | 0.006 | |
| B3 | DINOv2+ConvNeXt ZS | ViT-B/14 | Ensemble | No | 0.341 | |
| B2 | EfficientNet-B3 | EffNet-B3 | Single-view | No | 0.555 | 0.084 |
| B9 | VMamba+Mamba | VMamba-B | 2 Mamba | No | 0.717 | 0.052 |
| B6 | VMamba+Mamba | VMamba-B | 2 Mamba | Yes | 0.743 | 0.048 |
| E5 | DINOv3+FullMamba | DINOv3-L | 2 Mamba | No | 0.793 | 0.034 |
| E7 | DINOv3+4GDWC | DINOv3-L | 4 GDWC | No | 0.814 | 0.039 |
| E4 | DINOv3+Identity | DINOv3-L | Identity | No | 0.819 | 0.055 |
| E1 | DINOv3+BidirMamba | DINOv3-L | 2 BidirM | No | 0.819 | 0.051 |
| E6 | DINOv3+1GDWC | DINOv3-L | 1 GDWC | No | 0.821 | 0.034 |
| E8 | DINOv3+Identity+M | DINOv3-L | Identity | Yes | 0.828 | 0.053 |
| B8 | DINOv3+BidirM+M | DINOv3-L | 2 BidirM | Yes | 0.829 | 0.042 |
| E3 | DINOv3+GDWC+M | DINOv3-L | 2 GDWC | Yes | 0.829 | 0.045 |
| B7 | DINOv3+CVGA+M | DINOv3-L | 2 CVGA | Yes | 0.830 | 0.050 |
| E2 | DINOv3+CVGA | DINOv3-L | 2 CVGA | No | 0.833 | 0.051 |
| B4 | DINOv2+GDWC | DINOv2-L | 2 GDWC | No | 0.853 | 0.097 |
| B5 | DINOv3+GDWC | DINOv3-L | 2 GDWC | No | 0.903 | 0.064 |
IV-E Fusion Complexity Analysis
Table˜IV isolates fusion complexity effects on DINOv3-ViT-L without metadata.
| Rank | Fusion | ID | Std | Complexity | |
| 1 | 2 GatedDWConv | B5 | 0.903 | 0.064 | |
| 2 | 2 CVGA | E2 | 0.833 | 0.051 | |
| 3 | 1 GatedDWConv | E6 | 0.821 | 0.034 | |
| 4 | Identity (no fusion) | E4 | 0.819 | 0.055 | |
| 5 | 2 BidirMamba | E1 | 0.819 | 0.051 | |
| 6 | 4 GatedDWConv | E7 | 0.814 | 0.039 | |
| 7 | 2 FullMamba | E5 | 0.793 | 0.034 |
Three patterns emerge. (1) BidirMamba () matches the no-fusion identity baseline despite having the highest parameter count (17.55M), while FullMamba () falls 2.6 points below identity with 13.34M fusion parameters, establishing a negative correlation between fusion parameters and performance. (2) The GatedDWConv depth curve traces an inverted U: 0 blocks () 1 () 2 () 4 (). The disproportionate 1-to-2 block jump (+8.2 points) suggests that the 9-token receptive field captures a critical spatial scale at the left-right boundary, while the 2-to-4 drop (8.9) signals overfitting. (3) CVGA () outperforms both SSM variants but trails 2 GatedDWConv by 7.0 points: quadratic attention’s expressiveness is offset by overfitting on 286 training images per fold. Figure˜9 visualizes these results.
IV-F The Metadata Paradox
Table˜V presents a factorial crossing fusion type with metadata on/off on DINOv3-ViT-L.
| Fusion | No Meta | With Meta | |
| 2 GatedDWConv | 0.903 (B5) | 0.829 (E3) | 0.074 |
| 2 CVGA | 0.833 (E2) | 0.830 (B7) | 0.003 |
| 2 BidirMamba | 0.819 (E1) | 0.829 (B8) | 0.010 |
| Identity | 0.819 (E4) | 0.828 (E8) | 0.009 |
| (Fusion Effect) | 0.084 | 0.001 |
Without metadata, the fusion spread is 8.4 points, and with metadata, it collapses to 0.1 points as all configurations converge to . Metadata destroys the best model: GatedDWConv drops from to (7.4 points). The mechanism is straightforward: during training, species and state metadata provide predictive shortcuts through the MLP, and at test time, metadata is absent, and GatedDWConv suffers the largest degradation from this distribution shift. The column reveals an asymmetry: weaker models (Identity, BidirMamba) gain slightly ( to ), CVGA shows near zero effect (), and GatedDWConv is harmed most (). Metadata harm is proportional to model quality: the better the visual backbone, the more important it is to exclude training only metadata.
IV-G Backbone Pretraining Scale
Table˜VI controls for backbone scale with matched fusion and no metadata.
| Backbone | Pretrain Data | Params | ||
| EfficientNet-B3 | ImageNet-1K (1.2M) | 10.70M | 0.555 | baseline |
| VMamba-Base | ImageNet-1K (1.2M) | 88.56M | 0.717 | 0.162 |
| DINOv2-ViT-L | LVD-142M | 304.37M | 0.853 | 0.298 |
| DINOv3-ViT-L | LVD-1.7B | 303.08M | 0.903 | 0.348 |
Performance is strictly monotonic with pretraining scale. The DINOv2DINOv3 transition keeps architecture fixed while increasing pretraining data from 142M to 1.7B images, yielding points with zero additional parameters. Across the full backbone range, pretraining scale contributes points (EfficientNet-B3 to DINOv3), far exceeding the maximum fusion () or metadata () effects. VMamba-Base (), despite 8.3 more parameters than EfficientNet-B3, remains far below DINOv2 (), confirming that pretraining data dominates architecture choice. Practically, upgrading from DINOv2 to DINOv3 ( points, zero additional parameters, no overfitting risk) provides the most efficient single improvement, and the log-linear relationship in Figure˜8(b) suggests continued data scaling would yield predictable gains.
IV-H Feature Space Analysis
Figure˜10 examines the relationship between image derived color indices, sensor metadata, and biomass targets. The Excess Green Index (ExG) is moderately correlated with GDM () and Image Greenness with Dry Green (), whereas Mean Brightness shows negligible association with Dry Dead (), indicating that simple color features capture green biomass signals, but fail for senescent material. The NDVIHeight hexbin plot reveals that these two sensor metadata variables jointly stratify mean Dry Total across most of its range, explaining why metadata fusion improves weaker backbones, but saturates for DINOv3, which already encodes equivalent visual cues. The state level density plot further shows that geographic provenance induces distinct clusters in the NDVI–Height space, motivating the stratified group cross validation protocol adopted in Section˜IV. Figure˜11 shows spatial feature maps overlaid on representative images: DINOv3 produces spatially coherent activations that cleanly segment green vegetation from dead material, DINOv2 shows similar but less refined selectivity, and VMamba-Base produces coarser maps, paralleling the ranking and confirming that spatial discrimination, not fusion capacity, is the primary bottleneck.
IV-I Fold Analysis and Stability
Figure˜12 shows per-fold for all DINOv3-based models. Fold 4 is consistently hardest across every model (5–12 point drops), suggesting edge case pasture conditions: under represented species, atypical seasonal states, and a higher proportion of low biomass clover dominant swards, rather than a model specific failure. Table˜VII details Fold 4 performance: the proposed model (B5) achieves the highest score (0.779), but shows the largest drop (0.125), revealing a performance stability tradeoff.
| Model | Fold 4 | Drop |
| E4: Identity | 0.732 | 0.086 |
| E1: BidirMamba | 0.746 | 0.073 |
| B7: CVGA+Meta | 0.749 | 0.081 |
| E7: 4GDWC | 0.749 | 0.065 |
| E2: CVGA | 0.752 | 0.081 |
| E3: GDWC+Meta | 0.758 | 0.071 |
| B8: BidirMamba+Meta | 0.761 | 0.068 |
| E6: 1GDWC | 0.761 | 0.060 |
| B5: Proposed | 0.779 | 0.125 |
Table˜VIII presents stability analysis: the proposed model’s CV of 7.0% is moderate, while the most stable models (FullMamba 4.3%, 1GatedDWConv 4.2%) achieve lower . Metadata models cluster at lower CV (5.1–6.0%) because the ceiling suppresses both peaks and troughs, and practitioners valuing consistency may prefer 1 block GatedDWConv (CV4.2%, ).
| Model | Std | CV (%) | |
| E5: FullMamba | 0.793 | 0.034 | 4.3 |
| E6: 1GDWC | 0.821 | 0.034 | 4.2 |
| E7: 4GDWC | 0.814 | 0.039 | 4.7 |
| B8: BidirMamba+M | 0.829 | 0.042 | 5.1 |
| E3: GDWC+Meta | 0.829 | 0.045 | 5.4 |
| B7: CVGA+Meta | 0.830 | 0.050 | 6.0 |
| E1: BidirMamba | 0.819 | 0.051 | 6.2 |
| E2: CVGA | 0.833 | 0.051 | 6.2 |
| E4: Identity | 0.819 | 0.055 | 6.7 |
| B5: Proposed | 0.903 | 0.064 | 7.0 |
| B4: DINOv2+GDWC | 0.853 | 0.097 | 11.4 |
| B2: EffNet-B3 | 0.555 | 0.084 | 15.0 |
IV-J Prediction Quality Analysis
Figure˜13 shows predicted vs. actual scatter plots for all five targets () of the proposed model (B5, ). Predictions cluster tightly around the identity line, with larger residuals only for high-biomass tail samples, and Dry Clover shows the widest scatter, consistent with its 37.8% zero-inflation. Residuals are symmetric and centered near zero, confirming no systematic bias, and the tight Total scatter is encouraging given that compositional heads propagate component-level errors.
V Discussion
Implications for agricultural vision pipelines. Since global self attention is already performed across 24 transformer layers within each view by the DINOv3-ViT-L backbone, the fusion module need only enable local cross view communication, a role adequately served by kernel-5 depthwise convolution. Global mechanisms (attention transformers, SSMs) introduce parameters that overfit on images per fold, a scale typical of precision agriculture datasets. A general design principle is thus indicated: fusion complexity should be matched to dataset scale, not task aspiration.
Cautionary lessons for metadata fusion in agriculture. A broader risk for agricultural pipelines that fuse heterogeneous data sources is revealed by the metadata paradox: auxiliary features available only at training time can create harmful shortcuts. Predictive cues from species and state metadata (like “Lucerne in Victoria”) cause information to be routed through the metadata MLP at the expense of visual feature learning, and at test time, the strongest visual learner (GatedDWConv) suffers the largest degradation ( points). This finding is generalizable: any pipeline fusing training only auxiliary data (sensor readings, weather logs, field management records) risks the same shortcut, and modality dropout alone may be insufficient on small datasets.
Benchmarking position of CSIRO Biomass Dataset. Compared to GrassClover [33] (435 images, 2 Danish sites, May–October, no dead matter labels, and controlled lighting), CSIRO Biomass Dataset provides 2.7 greater geographic diversity (19 sites, 4 states), year round coverage, dead matter annotations, camera diversity, and heterogeneous auxiliary inputs (NDVI, and height) absent from any comparable benchmark. Other agricultural vision datasets (CropHarvest, PlantNet, DeepWeeds, Agriculture-Vision) target classification, detection, or aerial segmentation, and none provides ground level, component wise biomass regression with laboratory validated measurements. CSIRO Biomass Dataset is thus the only appropriate benchmark for this study, and the 17 configuration suite serves as a reproducible reference for future work.
Limitations. All findings are derived from a single dataset, though no comparable alternative with laboratory validated, component wise biomass ground truth for proximal pasture imagery currently exists. The fusion complexity inversion may not hold where global modules have sufficient data to avoid overfitting. The backbone comparison partially confounds pretraining data size with algorithmic improvements, and 8 GB VRAM limits batch sizes to 1. Validation on emerging agricultural benchmarks with similar ground truth quality is left to future work.
VI Conclusion
Foundation model adaptation for agricultural imagery is systematically studied on the 357 image CSIRO Pasture Biomass benchmark through 17 configurations. Three findings are established. First, fusion complexity inversion: a two layer gated depthwise convolution () outperforms cross view attention transformers (), bidirectional SSMs (), and full SSMs (, below the no fusion baseline). Second, foundation model dominance: rises monotonically from EfficientNet-B3 () to DINOv3-ViT-L (), confirming representation quality as the primary bottleneck. Third, the metadata fusion trap: training only metadata creates a ceiling at , collapsing an 8.4 point spread to 0.1 points. For scarce agricultural data, backbone quality should be prioritized, local fusion preferred, and inference-unavailable modalities excluded.